OpenStudy (anonymous):

Two dice are rolled. What is the probability that the product of the numbers rolled is even?

OpenStudy (raden):

hint : if all elements in the first dice is odd number then all elements the 2nd dice is even number to get the product of the numbers rolled is even for example : (1,2),(1,4),(1,6)... etc if all elements in the first dice is event number then can be order all elements the 2nd dice (odd and event number) to get the product of the numbers rolled is even for example : (2,1),(2,2),(2,3)... etc

OpenStudy (anonymous):

Wouldn't it be 1/4 + 1/4 + 1/4= 3/4?

OpenStudy (raden):

yeah, i got the answer 27/36 = 3/4

OpenStudy (anonymous):

Thanks!

OpenStudy (raden):

yw:)